OpenStudy (anonymous):

a refrigerator costs $560. the refrigerator is on sale for 30%off the regular price. what is the total cost of the refrigerator after the discount?

OpenStudy (anonymous):

560 multiply .30

OpenStudy (anonymous):

I thought u divide not multiply?#confused

OpenStudy (anonymous):

u multyply

OpenStudy (anonymous):

if u multy ply decimal it will go lower so the discout of the real price will go down

OpenStudy (anonymous):

oka. hoold im about to see what the answer is when u multiply.

OpenStudy (anonymous):

yeah but what is the discount?

OpenStudy (anonymous):

wats the answer then

OpenStudy (anonymous):

after u multiply or divide?

OpenStudy (anonymous):

the answer is 168.00

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

oh oka.. r u sure cuz i cant be getin my work wrong? im not tryin to say u r wrong i just want to know if its right or not?

OpenStudy (anonymous):

total price after discount is initial cost minus discount

OpenStudy (anonymous):

u dont minus

OpenStudy (anonymous):

wat grade are u in

OpenStudy (anonymous):

oka what im i suppose to do i have to people that r not sure bout what you do?

OpenStudy (anonymous):

go to the internet search it up

OpenStudy (anonymous):

the item is on sale for 30% less of it's initial price

OpenStudy (anonymous):

here r some choices a.)$728 b.)$175 c.)$392 d.)$168 is this better?

OpenStudy (anonymous):

168 is the discout that is (0.3* 560)

OpenStudy (anonymous):

please say you r sure bout this?!

OpenStudy (anonymous):

so final price is $392

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

I am 100% sure abt this

OpenStudy (anonymous):

how its 392

OpenStudy (anonymous):

Initial price-discount=final discount price

OpenStudy (mathteacher1729):

> a refrigerator costs $560. the refrigerator is on sale for 30%off the regular price. what is the total cost of the refrigerator after the discount? I'm guessing the regular price = $560 ? If that's the case, then there are two ways to figure it out, one of them not even requiring a calculator. Method one: Since you're lowering the price by 30%, that means that 70% of the original price will remain. So use your calculator and multiply 560 times 0.70. Method two: Since you're lowering the price by 30%, that means you're lowering it by 3 times 10%. 10 % of any number is easy to find (move the decimal to the left by one) 10% of 560 = 56 20% of 560 = 2*56 = 112 30% of 560 = 3*56 or 56+112 = 168 But we're SUBTRACTING 168 from 560. So what is 560 - 168? It's the same as 0.70 * 568. (check this with your calculator!)

OpenStudy (anonymous):